OpenStudy (anonymous):

I'm having trouble finding the equation of the line through the given points : (-1,8) & (3, -4). Thanks in advance!! :)

OpenStudy (anonymous):

First try to get the slope of the line, using the formula: \[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

I'm getting -3

OpenStudy (phi):

Did you get the equation? You use \[y-y_{1}= m(x-x_{1})\] and then use algebra to re-arrange into the form y = mx + b